miranda visited a petting zoo on saturday. she counted 95 ducks and goats.
she also counted 316 legs between the ducks and goats. how many of each kind of each animal did miranda see?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

miranda visited a petting zoo on saturday. she counted 95 ducks and goats.
she also counted 316 legs between the ducks and goats. how many of each kind of each animal did miranda see?
------------------------
Equations:
Heads: d + g = 95
Legs::2d +4g = 316
------------------------
Multiply thru the 1st equation by 2 to get:
2d + 2g = 190
------
Subtract that from the Legs equation and solve for "g":
2g = 126
g = 63 (# of goats)
---
Since d+g = 95, d = 32 (# of ducks)
----
Cheers,
Stan H.

Answer by mananth(16949)

let ducks be x numbers
goats be y numbers
..
x+y = 95....................(1)
..
2x+4y = 316.................(2)
multiply (1) by -2
-2x-2y=-190
add this to (2)
4y-2y=316-190
2y=126
/2
y=63 The number of goats.
plug value of y in (1)
x+63=95
x= 32 ... The number of ducks.
...
m.ananth@hotmail.ca